# Flying Bird Object Detection Algorithm in Surveillance Video

### v1

@@ IV Experiment
In this part, quantitative and qualitative experiments will be conducted to demonstrate the effectiveness of the proposed FBOD-SV. Next, we will describe the dataset (IV-A), evaluation method (IV-B), experimental platform (IV-C), implementation details (IV-D), ablation comparative analysis experiments (IV-E), cross-validation experiments (IV-F), and other methods comparative analysis experiments (IV-G).
@@ IV-B Evaluation Metrics
Same as [25], average precision(AP), the evaluation metric of Pascal VOC 2007 [22], is used to evaluate the detection results of the model. Specifically, AP50(The subscript 50 indicates that a test result is considered a true positive when the IOU between the test result and the true value is greater than or equal to 50%. That is, the IOU threshold is set to 50%), AP75(subscript 75 has a similar meaning to subscript 50), and AP(average accuracy exceeds multiple threshold averages, and the IOU threshold is set from 50% to 95% with 5% interval).
@@ IV-D Implementation Details
The feature aggregation unit of the model designed in this paper inputs n consecutive 3-channel RGB images of size 672 $\times$ 384, and outputs (n $\times$ 3 + 1) channel feature aggregation images of size 672 $\times$ 384. The input of the prediction network is the output of the feature aggregation unit (that is, a (n $\times$ 3+1) channel feature aggregation image of size 672 $\times$ 384), and the output is a 336 $\times$ 192 $\times$ 1 confidence prediction feature map and a 336 $\times$ 192 $\times$ 4 position regression feature map. The output predicts the position of the object on the intermediate frame.
@@
In this paper, all experiments are implemented under the Pytorch framework. All network models are trained on an NVIDIA GeForce RTX 3090 with 24 GB of GPU memory. For the batch size setting, the training model designed in this paper is set to 8, and other comparison experimental models are set according to their model size. All experimental models were trained from scratch without pre-trained models. The initial learning rate is set to 0.001, and the ratio decreases during training (per iteration, multiplied by 0.95). Meanwhile, based on [25], data augmentation adds enhancement operations such as random Cropping, random Center Flipping, and random HSV (All data augmentation performs the same operation on n consecutive frames, so Mosaic and Mixup data augments are inappropriate). The model is trained for a total of 100 iterations.
IV-E Ablation comparative analysis experiment
@@
In this subsection, three ablation contrast experiments will be set up according to the solution method of the challenges faced by detecting the flying bird object in the surveillance video to prove the effectiveness of the proposed solution. In particular, there are comparison experiments between single-frame image input and continuous multiple-frame image input, between multi-scale model structure and single-scale model structure, and between different label assignment strategies.
@@
IV-E1 Comparison experiment between single frame image input and continuous multiple frame image input
@@
In most cases, the characteristics of the flying bird in a single frame image are not obvious, and the features of the flying bird object cannot be extracted, so it is difficult to detect the flying bird object in the surveillance video. By aggregating the features of consecutive multi-frame images, the features of the flying bird object can be enriched so the flying bird object can be detected. The experimental results are shown in TABLE II. Compared with the single-frame image input, the $\text{AP}_{50}$ of continuous multi-frame image input (aggregation method based on ConvLSTM [25]) is increased by about 1.5 times, reaching 73.3%, which proves the above point. At the same time, compared with the ConvLSTM-based aggregation method [25]), the $\text{AP}_{50}$ of the aggregation method based on Co-Attention-FA is increased by 2.8%, which indicates that the feature aggregation method based on Co-Attention-FA proposed in this paper has more advantages.
TABLE II: Effect of continuous image input with different number of frames on detection performance.
Frames Aggregation Method $\text{AP}_{50}$ $\text{AP}_{75}$ AP
11 w/o 0.484 0.224 0.141
5 ConvLSTM 0.733 0.183 0.319
5 Co-Attention-FA 0.761 0.370 0.395
@@
    To be consistent with the conditions of other experiments, Mosaic and Mixup image enhancements were not applied to the single-frame image input. Therefore, its accuracy will be slightly lower than other advanced image-based object detection methods.
IV-E2 Comparison experiment between multi-scale model structure and single-scale model structure
@@
TABLE III shows the performance comparison between the multi-scale and single-scale models when it detects the flying bird object. It can be seen from the results that the detection accuracy of the single-scale model structure is even better than that of the multi-scale model structure ($\text{AP}_{50}$ is increased by 1.9%). Meanwhile, the inference speed of the single-scale model structure is slightly faster than that of the multi-scale model structure. Considering comprehensively, the single-scale model structure is more suitable for detecting flying birds in surveillance videos.
TABLE III: Effect of multi-scale model structure and single-scale model structure on detection performance.
Model Structure Frames $\text{AP}_{50}$ $\text{AP}_{75}$ AP Run time(s)
multi-scale 5 0.742 0.272 0.341 0.0072
single-scale 5 0.761 0.370 0.395 0.0069
IV-E3 Comparative experiments of different label assignment strategies
@@
TABLE IV shows the detection performance comparison results of the flying bird object detection models trained by the three label assignment strategies. As can be seen from the results, the detection accuracy $\text{AP}_{50}$ of the Center Gaussian label assignment method is 3.1% higher than that of the Shrinking Bounding Box label assignment method, which is because the Center Gaussian gives lower weight to the anchor point samples close to the boundary of the bounding box compared with the Shrinking Bounding Box label assignment method, thereby weakening the impact of inaccurate sample assignment. The detection accuracy $\text{AP}_{50}$ of the proposed SimOTA-OC label assignment method is 1.3% higher than that of the Center Gaussian label assignment method, which proves that the dynamic label assignment strategy is superior to the static label assignment strategy for the flying birds in the surveillance video.
TABLE IV: Effect of Diffrenet Label Assignment Method on detection performance.
@@
Shrinking Bounding Box 5 0.717 0.275 0.350
Center Gaussian 5 0.748 0.365 0.392
SimOTA-OC 5 0.7610 0.370 0.395
IV-F Cross-validation Experiments
To further verify the effectiveness of the proposed FBOD-SV, we add cross-validation experiments. Specifically, we divided the 120 videos into five groups of 24 videos each, as shown in Fig. 10. We set up five sets of experiments. Each set of experiments will alternately select one set of videos as the validation set. In contrast, the other four sets of videos will be used as the training set.
Figure 10: Video grouping situation.
The experimental results are shown in TABLE V. The experimental results show that (combined with the video grouping) when the scenarios in the validation Set all appear in the training set, the detection performance of the proposed method is better. For example, the $\text{AP}_{50}$ of Set 1 reaches 75.8%. Since there is at least one video scenario in Set 2 to Set 5 that is not in the training set, the detection performance of Set 2 to Set 5 shows varying degrees of decline, but its $\text{AP}_{50}$ is still high, which is enough to show that the bird object detection algorithm proposed in this paper is effective.
TABLE V: Algorithm performance when different video set is used as validation set.
Validation set $\text{AP}_{50}$ $\text{AP}_{75}$ AP
Set 1 as validation set 0.758 0.370 0.383
Set 2 as validation set 0.686 0.238 0.341
Set 3 as validation set 0.652 0.155 0.260
Set 4 as validation set 0.756 0.384 0.397
Set 5 as validation set 0.701 0.355 0.367
@@ IV-G Compared with other methods
The quantitative comparison experimental results are shown in TABLE VI. The $\text{AP}_{50}$ of the proposed FBOD-SV reaches 76.1%, which is 24.8% higher than that of SSD [10], 22.5% higher than that of YOLOV8l [19], 33.3% higher than that of SELSA [34], and 3% higher than FBOD-BMI [25].
@@
SSD [10] not only uses single-frame image as input, but its label allocation strategy is also static. However, the flying bird object in the surveillance video is not regular in the bounding box, so the static label allocation strategy will have more wrong assignments, affecting the algorithm’s accuracy. YOLOV8l [19] adopts the Task Aligned Assigner dynamic label assignment strategy and excellent model structure, so its detection performance is stronger than other object detection methods. However, it also uses single-frame image as input, and the effect is still not ideal when applied to the flying birds in the surveillance video because most of the single-frame image features of the flying birds in the surveillance video are not obvious. Video object detection methods use the information from multiple frames of images. However, they extract the information of a single frame of image in the initial feature extraction, which will lead to the loss of the bird object with unobvious features in the process of feature extraction and the formation of wrong features in the feature aggregation, resulting in the decline of detection accuracy. FBOD-BMI [25] adopts a two-stage method to make full use of the motion information of the flying bird object to detect the flying bird object. At the same time, ConvLSTM is used at the input of the model to aggregate the spatio-temporal information of adjacent multiple frames of the flying bird object, enhance the characteristics of the flying bird object, and improve the detection rate of the flying bird object. Therefore, FBOD-BMI [25] has achieved good results. Compared with FBOD-BMI [25], the proposed FBOD-SV uses Co-Attention-FA to aggregate the spatio-temporal information of flying bird objects. At the same time, a detection model more suitable for the flying bird object in the surveillance video is designed, the SimOTA-OC dynamic label allocation strategy is adopted, and the $\text{AP}_{50}$ of the proposed FBOD method is increased by 3%.
TABLE VI: Comparison with other object detection methods.
Method backbnoe $\text{AP}_{50}$ $\text{AP}_{75}$ AP $\text{AP}_{\text{S}}$ $\text{AP}_{\text{M}}$ $\text{AP}_{\text{L}}$
YOLOV5l [15] CSPResNet50 0.503 0.312 0.291 0.093 0.465 0.437
YOLOV6l [17] EfficientRep 0.519 0.320 0.219 0.079 0.481 0.643
YOLOV7l1 [18] CSPDarknet53 0.423 0.253 0.239 0.055 0.416 0.378
YOLOV8l [19] CSPResNet50 0.536 0.345 0.322 0.085 0.479 0.634
YOLOXl [16] CSPResNet50 0.528 0.349 0.306 0.089 0.491 0.596
SSD [10] SSDVGG16 0.513 0.283 0.275 0.073 0.449 0.496
Foveabox [43] ResNet101 0.434 0.128 0.188 0.033 0.338 0.418
CenterNet [50] ResNet18 0.469 0.281 0.256 0.056 0.428 0.519
FGFA [27] ResNet101 0.281 0.212 0.175 0.001 0.338 0.185
SELSA [34] ResNet101 0.428 0.290 0.248 0.017 0.471 0.565
Temporal RoI Align [35] ResNet101 0.424 0.279 0.247 0.018 0.469 0.519
FBOD-BMI2 [25] CSPDarkNet53 0.731 0.323 0.356 0.186 0.487 0.266
FBOD-SV (this paper) CSPDarkNet53 0.761 0.370 0.395 0.201 0.544 0.659
@@
In qualitative experiments, we selected three representative methods (YOLOV8 [19], SELSA [34], and FBOD-BMI [25]) for experiments and selected three typical scenarios to compare and analyze the experimental results. In scenario 1, the characteristics of the flying bird object are not obvious in the single-frame image, but the movement is obvious in the continuous image frames. In scenario 2, the appearance shape of the flying bird object changes greatly during its flight. In scenario 3, the characteristics of the single frame image of the flying bird object are not obvious, and the movement on the continuous image frames is not obvious either.
The experimental results are shown in Fig. 11. From the experimental results, it can be seen that in scenario 1, YOLOV8 [19] and SELSA [34] basically cannot detect the flying bird object, while FBOD-BMI [25] and FBOD-SV can detect the flying bird object well, which confirms the analysis in the quantitative experiment: that is, YOLOV8 [19] and SELSA [34] cannot effectively extract the features of flying bird objects when most single-frame image features of flying bird objects are not obvious. In scenario 2, the detection box of FBOD-BMI [25] is not as accurate as that of YOLOV8 [19] and FBOD-SV (the detection box of FBOD-BMI [25] in scenario 1 is not as accurate as that of FBOD-SV either), which indicates that the effect of static label allocation strategy is worse than that of dynamic label allocation strategy. In Scenario 3, all methods do not achieve good results. The reason for YOLOV8 [19] and SELSA [34] detection failure is the same as scenario 1. FBOD-BMI [25] and FBOD-SV aggregate the features of the flying bird object on consecutive frames of images. When the image features of the flying bird object are not obvious on consecutive frames, the detection will fail (In the subsequent research, we are considering using the global information of the video to solve this problem).
FBOD-SV FBOD-BMI  SELSA    YOLOV8l
(a) Scenario 1
(b) Scenario 2
(c) Scenario 3
Figure 11: Detection results of various object detection algorithms in three typical scenarios.

### v3

@@ IV Experiment
In this part, quantitative and qualitative experiments will be conducted to demonstrate the effectiveness of the proposed FBOD-SV. Next, we will describe the dataset (IV-A), evaluation method (IV-B), experimental platform (IV-C), implementation details (IV-D), parameter analysis experiment (IV-E), ablation comparative analysis experiments (IV-F), and other methods comparative analysis experiments (IV-G).
@@ IV-B Evaluation Metrics
Same as [26], average precision (AP), the evaluation metric of Pascal VOC 2007 [22], is used to evaluate the detection results of the model. Specifically, $\text{AP}_{50}$ (The subscript 50 indicates that a test result is considered a true positive when the IOU between the test result and the true value is greater than or equal to 50%. That is, the IOU threshold is set to 50%), $\text{AP}_{75}$ (subscript 75 has a similar meaning to subscript 50), and AP (average accuracy exceeds multiple threshold averages, and the IOU threshold is set from 50% to 95% with 5% interval).
@@ IV-D Implementation Details
Input $n$ consecutive 3-channel RGB images of size 672$\times$384 into the Co-Attention-FA, and output a feature aggregated image with ($n$$\times$3+1) channels and a size of 672$\times$384. The input of the FBOD-Net is the output of the Co-Attention-FA, and the output is a 336$\times$192$\times$1 confidence prediction feature map and a 336$\times$192$\times$4 position regression feature map. The output predicts the position of the object on the intermediate frame. Since only the object of the middle frame of consecutive $n$ frames is predicted when the input is a video, the first $(n-1)/2$ frames and the last $(n-1)/2$ frames have no detection result, as shown in Fig. 8LABEL:sub@Original ($n$=5 as an example). To solve this problem, this paper introduces video image sequence padding according to the principle of convolutional padding operation. A schematic diagram of the padding operation of the video image sequence is shown in Fig. 8LABEL:sub@After_padding, where an all-black image (each pixel of the image has a value of 0) is extended by $(n-1)/2$ frames at the beginning and end of the video image sequence. After the padding operation of the video image sequence, each frame of the original video sequence will have the prediction result.
Figure 8: Schematic diagram of input and output for detecting flying birds in a video using FBOD-SV. (a) No padding operation is applied to the video image sequence. (b) The padding operation is performed on the video image sequence.
@@
In this paper, all experiments are implemented under the Pytorch framework. All network models are trained on an NVIDIA GeForce RTX 3090 with 24 GB of GPU memory. For the batch size setting, the training model designed in this paper is set to 8, and other comparison experimental models are set according to their model size. All experimental models were trained from scratch without pre-trained models. The initial learning rate is set to 0.001, and the ratio decreases during training (per iteration, multiplied by 0.95). In addition, when training the model designed in this paper, based on [26], data augmentation adds enhancement operations such as random Cropping, random Center Flipping, and random HSV (All data augmentation performs the same operation on $n$ consecutive frames, so Mosaic and Mixup data augments are inappropriate). The model is trained for a total of 100 iterations.
IV-E Parameter Analysis Experiment
In this subsection, we conduct analysis experiments on the key parameters that affect the algorithm’s performance. Specifically, we analyze the effect of different numbers of consecutive input frames (the parameter $n$ mentioned above) on the algorithm’s performance. Three different numbers of consecutive input frames are set to test the influence on the algorithm’s performance. In particular, the effect of 3, 5, and 7 consecutive input frames on the algorithm’s performance is tested. In theory, the more consecutive input frames there are, the more information it can obtain. However, for the proposed method, as the number of consecutive input frames increases, the difficulty of information fusion will also increase, and the algorithm’s detection accuracy will not necessarily grow. In addition, the model inference time increases as the number of consecutive input frames increases.
The experimental results of the influence of different consecutive input frames on the algorithm’s performance are shown in TABLE II. The results show that when the number of consecutive input frames is 3, the algorithm’s accuracy is the lowest, but the running speed of the model is the fastest. The algorithm’s accuracy is the highest when the number of consecutive input frames is 5. When the number of consecutive input frames is 7, the running speed of the model is slow, and the algorithm’s accuracy is not the highest. Considering comprehensively, the number of consecutive input frames of the model is set to 5 in the process of subsequent experiments.
TABLE II: Effect of different numbers of consecutive input frames on the performance of the algorithm.
Frames $\text{AP}_{50}$ $\text{AP}_{75}$ AP Inference Time (s)
3 0.731 0.298 0.305 0.00697
5 0.762 0.371 0.395 0.00704
7 0.760 0.369 0.393 0.00712
IV-F Ablation Comparative Analysis Experiment
@@
In this subsection, three ablation contrast experiments will be set up according to the solution method of the challenges faced by detecting the flying bird object in the surveillance video to prove the effectiveness of the proposed solution. In particular, there are comparison experiments between single-frame image input and continuous multiple-frame image input IV-F1, between multi-scale model structure and single-scale model structure IV-F2, and between different label assignment strategies IV-F3.
@@
IV-F1 Comparison Experiment Between Single Frame Image Input and Continuous Multiple Frame Image input
@@
In most cases, the characteristics of the flying bird in a single frame image are not obvious, and the features of the flying bird object cannot be extracted, so it is difficult to detect the flying bird object in the surveillance video. By aggregating the features of consecutive multi-frame images, the features of the flying bird object can be enriched so the flying bird object can be detected. Fig. 9 shows the heatmaps of image features extracted by the models with different feature aggregation methods. It can be seen from the figure that after the feature aggregation of consecutive multiple frames of images, the features of the flying bird object, which are not obvious in the single-frame image, can be extracted.
(a) Raw Image
(b) w/o Aggregation
(c) ConvLSTM
(d) Co-Attention-FA
Figure 9: The Heatmaps of image features extracted by models using different feature aggregation methods. (b) Heatmap of image features extracted by the model without feature aggregation method (single frame image input). (c) Heatmap of image features extracted by the model based on ConvLSTM aggregation method (5 consecutive frames of image input). (d) Heatmap of image features extracted by the model based on the Co-Attention-FA aggregation method (5 consecutive frames of image input).
The experimental results of the quantitative comparison are shown in TABLE III. Compared with the single-frame image input, the $\text{AP}_{50}$ of continuous multi-frame image input (aggregation method based on ConvLSTM [26]) is increased by about 1.5 times, reaching 73.3%, which further supports the above view that continuous multi-frame feature aggregation can enhance the features of bird objects. Compared with the ConvLSTM-based aggregation method [26], the $\text{AP}_{50}$ of the aggregation method based on Co-Attention-FA is increased by 2.9%, and the model inference speed by 10.26%, which shows that the feature aggregation method based on Co-Attention-FA proposed in this paper is more advantageous.
TABLE III: Effect of different feature aggregation methods on algorithm performance.
Frames Aggregation Method $\text{AP}_{50}$ $\text{AP}_{75}$ AP Inference Time (s)
11 w/o 0.484 0.224 0.141 0.0069
5 ConvLSTM 0.733 0.184 0.319 0.0078
5 Co-Attention-FA 0.762 0.371 0.395 0.0070
@@
    To maintain consistency with the conditions of the other two comparison items, Mosaic and Mixup image enhancements were not applied to the single-frame image input. Therefore, its accuracy will be slightly lower than other advanced image-based object detection methods.
IV-F2 Comparison Experiment Between Multi-scale Model Structure and Single-scale Model Structure
@@
TABLE IV shows the performance comparison between the multi-scale and single-scale models when it detects the flying bird object. It can be seen from the results that the detection accuracy of the single-scale model structure is even better than that of the multi-scale model structure ($\text{AP}_{50}$ is increased by 1.8%). Meanwhile, the inference speed of the single-scale model structure is slightly faster than that of the multi-scale model structure. Considering comprehensively, the single-scale model structure is more suitable for detecting flying birds in surveillance videos.
TABLE IV: Effect of multi-scale model structure and single-scale model structure on detection performance.
Model Structure Frames $\text{AP}_{50}$ $\text{AP}_{75}$ AP Inference Time (s)
Multi-scale 5 0.744 0.276 0.343 0.0072
Single-scale 5 0.762 0.371 0.395 0.0070
IV-F3 Comparative Experiments of Different Label Assignment Strategies
@@
TABLE V shows the detection performance comparison results of the flying bird object detection models trained by the three label assignment strategies. As can be seen from the results, the detection accuracy $\text{AP}_{50}$ of the Center Gaussian label assignment method is 3.4% higher than that of the Shrinking Bounding Box label assignment method, which is because the Center Gaussian gives lower weight to the anchor point samples close to the boundary of the bounding box compared with the Shrinking Bounding Box label assignment method, thereby weakening the impact of inaccurate sample assignment. The detection accuracy $\text{AP}_{50}$ of the proposed SimOTA-OC label assignment method is 1.2% higher than that of the Center Gaussian label assignment method, which proves that the dynamic label assignment strategy is superior to the static label assignment strategy for the flying birds in the surveillance video.
TABLE V: Effect of different label assignment method on detection performance.
@@
Shrinking Bounding Box 5 0.716 0.276 0.349
Center Gaussian 5 0.750 0.370 0.393
SimOTA-OC 5 0.762 0.371 0.395
@@ IV-G Compared with Other Methods
The quantitative comparison experimental results are shown in TABLE VI. The $\text{AP}_{50}$ of the proposed FBOD-SV reaches 76.2%, which is 24.9% higher than that of SSD [10], 22.6% higher than that of YOLOV8l [19], 33.4% higher than that of SELSA [31], and 3.1% higher than FBOD-BMI [26]. At the same time, the detection speed of the proposed method is twice that of FBOD-BMI, reaching 59.87fps to meet the real-time requirement of the bird object detection task in surveillance video.
@@
SSD [10] not only uses single-frame image as input, but its label allocation strategy is also static. However, the flying bird object in the surveillance video is not regular in the bounding box, so the static label allocation strategy will have more wrong assignments, affecting the algorithm’s accuracy. YOLOV8l [19] adopts the Task Aligned Assigner dynamic label assignment strategy and excellent model structure, so its detection performance is stronger than other object detection methods. However, it also uses single-frame image as input, and the effect is still not ideal when applied to the flying birds in the surveillance video because most of the single-frame image features of the flying birds in the surveillance video are not obvious. Video object detection methods use the information from multiple frames of images. However, they extract the information of a single frame of image in the initial feature extraction, which will lead to the loss of the bird object with unobvious features in the process of feature extraction and the formation of wrong features in the feature aggregation, resulting in the decline of detection accuracy. FBOD-BMI [26] adopts a two-stage method to make full use of the motion information of the flying bird object to detect the flying bird object. At the same time, ConvLSTM is used at the input of the model to aggregate the spatio-temporal information of adjacent multiple frames of the flying bird object, enhance the characteristics of the flying bird object, and improve the detection rate of the flying bird object. Therefore, FBOD-BMI [26] has achieved good results. Compared with FBOD-BMI [26], the proposed FBOD-SV uses Co-Attention-FA to aggregate the spatio-temporal information of flying bird objects. At the same time, a detection model more suitable for the flying bird object in the surveillance video is designed, the SimOTA-OC dynamic label allocation strategy is adopted, and the $\text{AP}_{50}$ of the proposed FBOD method is increased by 3.1%.
[…]
Method Backbnoe Image Size $\text{AP}_{50}$ $\text{AP}_{75}$ AP $\text{AP}_{\text{S}}$ $\text{AP}_{\text{M}}$ $\text{AP}_{\text{L}}$ Detection Speed (fps)
YOLOV5l [15] CSPResNet50 640$\times$640 0.503 0.312 0.291 0.093 0.465 0.437 83.68
YOLOV6l [17] EfficientRep 640$\times$640 0.519 0.320 0.219 0.079 0.481 0.643 65.77
YOLOV7l1 [18] CSPDarkNet53 640$\times$640 0.423 0.253 0.239 0.055 0.416 0.378 78.61
YOLOV8l [19] CSPResNet50 640$\times$640 0.536 0.345 0.322 0.085 0.479 0.634 67.52
YOLOXl [16] CSPResNet50 640$\times$640 0.528 0.349 0.306 0.089 0.491 0.596 48.29
SSD [10] SSDVGG16 640$\times$640 0.513 0.283 0.275 0.073 0.449 0.496 59.12
Foveabox [36] ResNet101 640$\times$640 0.434 0.128 0.188 0.033 0.338 0.418 42.82
CenterNet [43] ResNet18 640$\times$640 0.469 0.281 0.256 0.056 0.428 0.519 38.44
FGFA [28] ResNet101 1000$\times$600 0.281 0.212 0.175 0.001 0.338 0.185 17.65
SELSA [31] ResNet101 1000$\times$600 0.428 0.290 0.248 0.017 0.471 0.565 14.88
Temporal RoI Align [32] ResNet101 1000$\times$600 0.424 0.279 0.247 0.018 0.469 0.519 7.28
FBOD-BMI2 [26] CSPDarkNet53 672$\times$384 0.731 0.323 0.356 0.186 0.487 0.266 28.56
FBOD-SV (this paper) CSPDarkNet53 672$\times$384 0.762 0.371 0.395 0.203 0.545 0.659 59.87
@@
In qualitative experiments, we selected three representative methods (YOLOV8l [19], SELSA [31], and FBOD-BMI [26]) for experiments and selected four typical Situations to compare and analyze the experimental results. The four Situations are shown in Fig. 11. In Situation 1, there are two flying bird objects with clear appearance features. In Situation 2, the characteristics of the flying bird object are not obvious in the single-frame image, but the movement is obvious in the continuous image frames. In Situation 3, the appearance shape of the flying bird object changes greatly during its flight. In Situation 4, the characteristics of the single frame image of the flying bird object are not obvious, and the movement on the continuous image frames is not obvious either.
(a) Situation 1
(b) Situation 2
(c) Situation 3
(d) Situation 4
Figure 11: Four typical Situations.
The experimental results are shown in Fig. 12. From the experimental results, it can be seen that in Situation 1, all methods can detect the flying bird objects (when the bird object is slightly similar to the background, YOLOV8l [19] and SELSA [31] will miss detection in the third and fourth frames). In Situation 2, YOLOV8l [19] and SELSA [31] basically cannot detect the flying bird object, while FBOD-BMI [26] and FBOD-SV can detect the flying bird object well, which confirms the analysis in the quantitative experiment: that is, YOLOV8l [19] and SELSA [31] cannot effectively extract the features of flying bird objects when most single-frame image features of flying bird objects are not obvious. In Situation 3, the detection box of FBOD-BMI [26] is not as accurate as that of YOLOV8l [19] and FBOD-SV (the detection box of FBOD-BMI [26] in Situation 1 is not as accurate as that of FBOD-SV either), which indicates that the effect of static label allocation strategy is worse than that of dynamic label allocation strategy. In Situation 4, all methods do not achieve good results. The reason for YOLOV8l [19] and SELSA [31] detection failure is the same as Situation 1. FBOD-BMI [26] and FBOD-SV aggregate the features of the flying bird object on consecutive frames of images. When the image features of the flying bird object are not obvious on consecutive frames, the detection will fail (In the subsequent research, we are considering using the global information of the video to solve this problem).
$1^{\text{{st}}}$ frame            $2^{\text{{nd}}}$ frame            $3^{\text{{rd}}}$ frame            $4^{\text{{th}}}$ frame
(a) Situation 1
$1^{\text{{st}}}$ frame            $2^{\text{{nd}}}$ frame            $3^{\text{{rd}}}$ frame            $4^{\text{{th}}}$ frame
(b) Situation 2
$1^{\text{{st}}}$ frame            $2^{\text{{nd}}}$ frame            $3^{\text{{rd}}}$ frame            $4^{\text{{th}}}$ frame
(c) Situation 3
$1^{\text{{st}}}$ frame            $2^{\text{{nd}}}$ frame            $3^{\text{{rd}}}$ frame            $4^{\text{{th}}}$ frame
(d) Situation 4
Figure 12: Screenshots of detection results on four consecutive frames by different methods in four typical Situations. The first row shows YOLOV8l detection results, the second row shows SELSA detection results, the third row shows FBOD-BMI detection results, and the fourth row shows FBOD-SV detection results.